# Ore extensions of multiplier Hopf coquasigroups

### v1

@@ 3 Ore extensions of multiplier Hopf coquasigroups
for any $a\in M(A).$
[…]
In fact,
$\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}$ $=\tau\Big{(}S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}\Big{)}=\tau\big{(}\chi({a_{(1)}})S({a_{(2)}})\big{)}=\chi({a_{(1)}})\tau\big{(}S({a_{(2)}})\big{)}$
$=\chi({a_{(1)}})A{d_{r}}\Big{(}S(a_{(2)(2)})\chi\big{(}S(a_{(2)(1)})\big{)}\Big{)}$
$=\chi(a_{(1)})S(a_{(2)(1)})Ad_{r}\big{(}S(a_{(2)(2)})\big{)}$
$=Ad_{r}\big{(}S(a)\big{)}$
[…]
$S\big{(}\tau(a)\big{)}=S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}=\chi(a_{(1)})S(a_{(2)}).$
[…]
$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=rS\big{(}\delta(a)\big{)}.$ (3.14)
[…]
Applying $m(\iota\otimes S)$ to the equation
@@
By applying the properties of the antipode, we have
@@
And considering the nondegeneracy of the product of $A$, this means that the following equation holds for any $a\in A$
@@
Hence, we have
$rS\big{(}\delta(a)\big{)}$ $=rS(a_{(1)})a_{(2)(1)}S\big{(}\delta(a_{(2)(2)})\big{)}$
=(3.15)${=}-rS(a_{(1)})r^{-1}\delta(a_{(2)(1)})S(a_{(2)(2)})$ )
$=-Ad_{r}S({{a_{(1)}}})\delta(a_{(2)(1)})S(a_{(2)(2)}),$ (3.16)
for any $a\in A$.
In addition, from the action of $\delta$ on the element $\varepsilon(a)=a_{1}S(a_{(2)})$, it follows that
@@
[…]
$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}$ $=\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})a_{(2)(2)(1)}\big{)}\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
@@ 4 Isomorphism of Ore extensions
From Theorem 3.4, it follows that a MHC-Ore extension is determined by $\chi$, $r$, and $\delta$. Thus, $A(\chi,r,\delta)$ is used to denote the MHC-Ore extension of $A[y;\tau,\delta]$, where $\chi$ is a character of $A$, $r$ is a group-like element in $M(A)$ and $\delta$ is a $\langle\chi,r\rangle$-derivation. The corresponding Ore extension on the multiplier coquasigroup is denoted by $M(A)(\chi,r,\delta)$.
@@
Similarly to [14, Theorem 4.9], we have the following theorem for regular multiplier Hopf coquasigroups.
Theorem 4.2. For multiplier Hopf coquasigroups $A$ and $A^{\prime}$, let their Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ be defined with coproducts: $\Delta(y)=y\otimes 1+r\otimes y$, $\Delta(y^{\prime})=y^{\prime}\otimes 1+r^{\prime}\otimes y^{\prime}$, where $r$ and $r^{\prime}$ are group-like elements in $A$. Then $A(\chi,r,\delta)$ is isomorphic to $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ if there is an isomorphism $\phi:A\rightarrow A^{\prime}$ such that
$\phi(r)=r^{\prime},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},$ (4.1)
@@
Proof. Let $\phi(y)=y^{\prime}+d^{\prime}$. Then $\phi$ can be extended from $\phi:A\rightarrow A^{\prime}$ to $\phi:A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$. We have
@@
$=\phi\big{(}\tau(a)\big{)}y^{\prime}+\phi\big{(}\tau(a)\big{)}d^{\prime}+\phi\big{(}\delta(a)\big{)}=\phi\big{(}\tau(a)\big{)}(y^{\prime}+d^{\prime})+\phi\big{(}\delta(a)\big{)},$
for any $a\in A$. Thus we get
@@
Considering the properties of $\varepsilon$, we obtain
$d^{\prime}=(\varepsilon\otimes id)\big{(}\Delta(d^{\prime})\big{)}=\varepsilon(d^{\prime})+\varepsilon(r^{\prime})d^{\prime}=\varepsilon(d^{\prime})+d^{\prime},$
this indicates $\varepsilon(d^{\prime})=0$. And by Proposition 3.3, we know that $\varepsilon(y)=0$ and $\varepsilon(y^{\prime})=0$. So we have
@@
this means that $\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y)$.
Considering the properties of $S$, we obtain
$0=\varepsilon(d^{\prime})=m(S\otimes id)\Delta(d^{\prime})=S(d^{\prime})+S(r^{\prime})d^{\prime}=S(d^{\prime})+r{{}^{\prime-1}}d^{\prime},$
this indicates $S(d^{\prime})=-r{{}^{\prime-1}}d^{\prime}$. Thus we have
$\phi\big{(}S(y)\big{)}=\phi(-{r^{-1}}y)=-\phi({r^{-1}})\phi(y)=-r{{}^{\prime-1}}(y^{\prime}+d^{\prime}),$
$S\big{(}\phi(y)\big{)}=S(y^{\prime}+d^{\prime})=S(y^{\prime})+S(d^{\prime})=-{r^{-1}}y^{\prime}-r{{}^{\prime-1}}d^{\prime}=-r{{}^{\prime-1}}(y^{\prime}+d^{\prime}),$

### v2

@@ 3 Ore extensions of multiplier Hopf coquasigroups
It remains to show that $(\ref{3.12})$ and $(\ref{3.13})$ hold for every $a\in A$. Indeed,
$\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}$ $\overset{(\ref{3.2})}{=}\tau\Big{(}S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}\Big{)}=\tau\big{(}\chi({a_{(1)}})S({a_{(2)}})\big{)}=\chi({a_{(1)}})\tau\big{(}S({a_{(2)}})\big{)}$
$\overset{(\ref{3.2})(\ref{3.3})}{=}\chi({a_{(1)}})A{d_{r}}\Big{(}S(a_{(2)(2)})\chi\big{(}S(a_{(2)(1)})\big{)}\Big{)}$
$=\chi({a_{(1)}})\chi\big{(}S(a_{(2)(1)})\big{)}A{d_{r}}\big{(}S(a_{(2)(2)})\big{)}$
$=\chi\big{(}a_{(1)}S(a_{(2)(1)})\big{)}Ad_{r}\big{(}S(a_{(2)(2)})\big{)}$
$\overset{(\ref{2.1})}{=}Ad_{r}\big{(}S(a)\big{)},$
which confirms $(\ref{3.12})$.
Proceeding to equation $(\ref{3.13})$, we utilize equation $(\ref{3.2})$, from which it follows that
$S\big{(}\tau(a)\big{)}\overset{(\ref{3.2})}{=}S\big{(}\chi({a_{(1)}}){a_{(2)}}\big{)}=\chi(a_{(1)})S(a_{(2)}).$
Thus, equation $(\ref{3.13})$ can be reformulated equivalently as
$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}=rS\big{(}\delta(a)\big{)},$ (3.14)
it suffices to demonstrate the validity of $(\ref{3.14})$. Applying the map $m(\iota\otimes S)$ to both sides of the identity
@@
Using the defining property of the antipode, the left-hand side simplifies as
@@
By the nondegeneracy of the multiplication in $A$, we conclude that the following identity holds for all $a\in A$
@@
Therefore, we have
$rS\big{(}\delta(a)\big{)}$ $\overset{(\ref{2.1})}{=}rS(a_{(1)})a_{(2)(1)}S\big{(}\delta(a_{(2)(2)})\big{)}$
$\overset{(\ref{3.15})}{=}-rS(a_{(1)})r^{-1}\delta(a_{(2)(1)})S(a_{(2)(2)})$
$=-Ad_{r}\big{(}S(a_{(1)})\big{)}\delta(a_{(2)(1)})S(a_{(2)(2)})$ (3.16)
for all $a\in A$. Furthermore, applying $\delta$ to the identity $\varepsilon(a)=a_{(1)}S(a_{(2)})$, we find
@@
By combining $(\ref{3.16})$ and $(\ref{3.17})$, we obtain the following
$\chi(a_{(1)})\delta\big{(}S(a_{(2)})\big{)}$ $\overset{(\ref{2.1})}{=}\chi(a_{(1)})\tau\big{(}S(a_{(2)(1)})a_{(2)(2)(1)}\big{)}\delta\big{(}S(a_{(2)(2)(2)})\big{)}$
@@
Remark 3.5. The proof of Theorem 3.4 adopts a strategy analogous to that of Theorem 3.3 in [3] and the theorem in [16], albeit with notable differences in technical execution. If the underlying algebra of $A$ is unital, then $A$ becomes a Hopf coquasigroup, recovering the framework studied in [3]. And if the comultiplication in $A$ is coassociative, then $A$ reduces to a multiplier Hopf algebra, corresponding to the setting treated in [16].
Example 3.6. $(1)$ Let $(A,\Delta,\varepsilon_{A},S_{A})$ be a multiplier Hopf algebra and $(H,\Delta,\varepsilon_{H},S_{H})$ a Hopf coquasigroup. Then the tensor product $A\otimes H$ forms a multiplier Hopf coquasigroup under the following structure:
$(a\otimes h)(a^{\prime}\otimes h^{\prime})=aa^{\prime}\otimes hh^{\prime},$
$\Delta(a\otimes h)=(a_{(1)}\otimes h_{(1)})\otimes(a_{(2)}\otimes h_{(2)})\in M\big{(}(A\otimes H)\otimes(A\otimes H)\big{)},$
$\varepsilon_{A\otimes H}(a\otimes h)=\varepsilon_{A}(a)\varepsilon_{H}(h),$
$S_{A\otimes H}(a\otimes h)=S_{A}(a)\otimes S_{H}(h),$
where $a,a^{\prime}\in A$ and $h,h^{\prime}\in H$. If $A[y;\tau,\delta]$ is an Ore extension of $A$, then $A[y;\tau,\delta]\otimes H$ is an Ore extension of the multiplier Hopf coquasigroup $A\otimes H$ with
$\Delta_{A\otimes H}(y\otimes 1_{H})=(y\otimes 1_{H})\otimes(1_{M(A)}\otimes 1_{H})+(r\otimes 1_{H})\otimes(y\otimes 1_{H}).$
Similarly, $A\otimes H[y;\tau,\delta]$ is an Ore extension of $A\otimes H$ with $\Delta_{A\otimes H}(1_{M(A)}\otimes y)=(1_{M(A)}\otimes y)\otimes(1_{M(A)}\otimes 1_{H})+(1_{M(A)}\otimes r)\otimes(1_{M(A)}\otimes y)$.
$(2)$ Let $A$ be an algebra spanned by elements $\{e_{p}e_{\alpha}\mid p\in\mathbb{Z},\alpha\in\mathbb{G}\}$ over $\mathbb{C}$, where $\mathbb{G}$ is a (finite or infinite) quasigroup, and the generators $e_{p}e_{\alpha}$ satisfy the orthogonality relations $(e_{p}e_{\alpha})(e_{q}e_{\beta})=\delta^{K}_{p,q}\delta^{K}_{\alpha,\beta}e_{p}e_{\alpha}$, with $\delta^{K}$ denoting the Kronecker delta.
Define a comultiplication $\Delta$ on $A$ by
$\Delta(e_{p}e_{\alpha})=\sum\limits_{\beta\in\mathbb{G}}\sum\limits_{k\in\mathbb{Z}}e_{k}e_{\beta}\otimes e_{p-k}e_{\beta^{-1}\alpha}.$
Note that these infinite sums are well-defined in the strict topology on the multiplier algebra, i.e., multiplication with elements of $A$ yields finite sums. One may verify that $(A,\Delta)$ is a regular multiplier Hopf coquasigroup. The counit is given by $\varepsilon(e_{p}e_{\alpha})=\delta^{K}_{p,0}\delta^{K}_{\alpha,e}$, and the antipode by $S(e_{p}e_{\alpha})=e_{-p}e_{\alpha^{-1}}$.
Following the construction in Example 3.9 of [16], choose a nonzero scalar $\lambda\in\mathbb{C}\setminus\{0\}$. Define $r\in M(A)$ as $r=\sum\limits_{k\in\mathbb{Z}}\lambda^{k}e_{k}$, a character $\chi:A\rightarrow\mathbb{C}$ by $\chi(e_{p}e_{\alpha})=\delta^{K}_{p+2,0}\delta^{K}_{\alpha,e}$, an algebra endomorphism $\tau(e_{p}e_{\alpha})=e_{p+2}e_{\alpha}$ and a $\tau$-derivation $\delta:A\rightarrow A$ by $\delta(e_{p}e_{\alpha})=0$. It is straightforward to verify that these definitions satisfy conditions (3.2)-(3.4). Consequently, the Ore extension $A[y;\tau,\delta]$ is established, thereby endowing AA with the structure of a multiplier Hopf coquasigroup in the sense of Theorem 3.4.
At the end of this section, we introduce the following theorem concerning the *-structure of MHC-Ore extensions. In parallel with Definition 2.4 presented in [10], we first define the notion of a multiplier Hopf *-coquasigroup.
Definition 3.7. If $A$ is a *-algebra over $\mathds{C}$, we call $\Delta$ a comultiplication if it is also a *-homomorphism. A multiplier Hopf *-coquasigroup is a *-algebra with a comultiplication, making it into a multiplier Hopf coquasigroup.
Parallel to the proposition in [10], we have the following proposition for multiplier Hopf *-coquasigroups.
Proposition 3.8. If $A$ is a multiplier Hopf *-coquasigroup, then for all $a\in A$, we have
$(1)$ $\varepsilon(a^{*})=\varepsilon(a)^{-}$ (where $\varepsilon(a)^{-}$ denotes the conjugate of $\varepsilon(a)$);
$(2)$ $S(S(a)^{*})^{*}=a$.
Proof. $(1)$ Define $\tilde{\varepsilon}(a)=\varepsilon(a^{*})^{-}$. Consider the action of $(\tilde{\varepsilon}\otimes\iota)$ on $(1\otimes a)\Delta(b)$
$(\tilde{\varepsilon}\otimes\iota)\big{(}(1\otimes a)\Delta(b)\big{)}$ $=\Big{(}(\varepsilon\otimes\iota)\big{(}\Delta(b^{*})(1\otimes a^{*})\big{)}\Big{)}^{*}=(b^{*}a^{*})^{*}=ab$
$=(\varepsilon\otimes\iota)\big{(}(1\otimes a)\Delta(b)\big{)}.$
This implies $\tilde{\varepsilon}=\varepsilon$. Substituting the definition of $\varepsilon$, we conclude that $\varepsilon(a)=\varepsilon(a^{*})^{-}$, and hence $\varepsilon(a^{*})=\varepsilon(a)^{-}$.
$(2)$ Suppose $a\otimes b=\sum_{i=1}^{n}\Delta(a_{i})(1\otimes b_{i})$. Applying $m(S\otimes\iota)$ to both sides yields
$S(a)b$ $=m(S\otimes\iota)(a\otimes b)=\sum_{i=1}^{n}m(S\otimes\iota)\big{(}\Delta(a_{i})(1\otimes b_{i})\big{)}$
$=\sum_{i=1}^{n}\big{(}m(S\otimes\iota)\Delta(a_{i})\big{)}b_{i}=\sum_{i=1}^{n}\varepsilon(a_{i})b_{i}.$
Thus, $S(a)b=\sum_{i=1}^{n}\varepsilon(a_{i})b_{i}$. Taking adjoints and applying the flip map gives $b^{*}\otimes a^{*}=\sum_{i=1}^{n}(b_{i}\otimes 1)\Delta^{cop}(a_{i}^{*})$. Next, applying $(\iota\otimes S^{-1})$ and multiplying on the right by $1\otimes c^{*}$, it follows from Lemma 3.1 in [15] that
$b^{*}S^{-1}(a^{*})=\sum_{i=1}^{n}b_{i}^{*}\varepsilon(a_{i}^{*})c^{*}.$
[…]
Theorem 3.9. Let $A$ be a multiplier Hopf *-coquasigroup. Then the MHC-Ore extension $A[y;\tau,\delta]$ is also a multiplier Hopf *-coquasigroup provided the following conditions are satisfied:
$(1)$ $\tau$ is an involution;
$(2)$ $\delta(\tau+\iota)=0$ and $\delta*=*\delta$;
$(3)$ $r^{*}=r$.
Proof. We begin by extending the *-operator to $M(A[y;\tau,\delta])$ such that $y^{*}=y$. For $y\in M(A)[y;\tau,\delta]$ and any $a\in A$, conditions (1) and (2) imply
$(ya)^{*}$ $=\big{(}\tau(a)y+\delta(a)\big{)}^{*}=y\tau(a)^{*}+\delta(a)^{*}$
$=\tau\big{(}\tau(a)^{*}\big{)}y+\delta\big{(}\tau(a)^{*}\big{)}+\delta(a)^{*}=a^{*}y+\delta\big{(}\tau(a)\big{)}^{*}+\delta(a)^{*}$
$=a^{*}y+\delta\big{(}\tau(a)+a\big{)}^{*}=a^{*}y+\big{(}\delta(\tau+\iota)(a)\big{)}^{*}$
$=a^{*}y=a^{*}y^{*}.$
Applying the result $a^{*}y=y\tau(a)^{*}+\delta(a)^{*}$ derived above, we have for any $a\in A$ and $x\in M(A)$
$a^{*}x^{*}y$ $=(xa)^{*}y=y\tau(xa)^{*}+\delta(xa)^{*}$
$=y(\tau(x)\tau(a))^{*}+(\delta(x)a+\tau(x)\delta(a))^{*}$
$=y\tau(a)^{*}\tau(x)^{*}+a^{*}\delta(x)^{*}+\delta(a)^{*}\tau(x)^{*}$
$=(y\tau(a)^{*}+\delta(a)^{*})\tau(x)^{*}+a^{*}\delta(x)^{*}$
$=(a^{*}y)\tau(x)^{*}+a^{*}\delta(x)^{*}$
$=a^{*}(y\tau(x)^{*}+\delta(x)^{*}),$
this implies $x^{*}y^{*}=y^{*}\tau(x)^{*}+\delta(x)^{*}$. Hence, the involution * can be extended to a map $M(A)\to M(A)$ such that $y^{*}=y$. Moreover, for any element $\sum_{i=1}^{n}a_{i}y^{i}\in A[y;\tau,\delta]$, its image under the involution satisfies $(\sum_{i=1}^{n}a_{i}y^{i})^{*}\in A[y;\tau,\delta]$. It is straightforward to verify that $A[y;\tau,\delta]$ forms a *-algebra under the above definition.
Next, we prove that $\Delta$ is a *-homomorphism. From condition (3), it follows that
$\Delta(y)^{*}$ $=(y\otimes 1+r\otimes y)^{*}=y^{*}\otimes 1+r^{*}\otimes y^{*}$
$=y\otimes 1+r\otimes y=\Delta(y)$
$=\Delta(y^{*}).$
Since $\Delta$ is an algebra homomorphism, it implies $\Delta(\sum_{i=1}^{n}a_{i}y^{i})^{*}=\Delta\big{(}(\sum_{i=1}^{n}a_{i}y^{i})^{*}\big{)}$.
Furthermore, it is clear that $\varepsilon(y^{*})=0=\overline{\varepsilon(y)}$, and more generally, $\varepsilon\big{(}(\sum_{i=1}^{n}a_{i}y^{i})^{*}\big{)}=\overline{\varepsilon(\sum_{i=1}^{n}a_{i}y^{i})}$, so $\varepsilon$ is a *-homomorphism.
It remains to prove that $S\big{(}S(y)^{*}\big{)}^{*}=y$. Using Proposition 3.3 and condition (3), we obtain
$S(y)^{*}$ $=(-r^{-1}y)^{*}=-y^{*}(r^{-1})^{*}$
$=-yr^{-1}=S(y).$
and therefore
$S\big{(}S(y)^{*}\big{)}$ $=S\big{(}S(y)\big{)}=S(-r^{-1}y)$
$=-S(y)S(r^{-1})=r^{-1}yr$
$=y.$
@@ 4 Isomorphism of Ore extensions
According to Theorem 3.4, an Ore extension of a regular multiplier Hopf coquasigroup is uniquely determined by the parameters $\chi$, $r$, and $\delta$. We therefore denote such an Ore extension as $A[y;\tau,\delta]$ by $A(\chi,r,\delta)$, where $\chi$ is a character on $A$, $r$ is a group-like element in $M(A)$, and $\delta$ is a $\langle\chi,r\rangle$-derivation. The corresponding Ore extension at the level of multiplier coquasigroups is denoted by $M(A)(\chi,r,\delta)$.
@@
Moreover, in the multiplier Hopf *-coquasigroup case, $A$ and $A^{\prime}$ are said to be isomorphic if furthermore $\phi$ is a *-homomorphism.
In parallel with [16, Theorem 4.9], we have the following theorem for regular multiplier Hopf coquasigroups.
Theorem 4.2. For multiplier Hopf coquasigroups $A$ and $A^{\prime}$, let their Ore extensions $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ be defined with communications: $\Delta(y)=y\otimes 1+r\otimes y$, $\Delta(y^{\prime})=y^{\prime}\otimes 1+r^{\prime}\otimes y^{\prime}$, where $r$ and $r^{\prime}$ are group-like elements in $A$. Then $A(\chi,r,\delta)$ is isomorphic to $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ if there is an isomorphism $\phi:A\rightarrow A^{\prime}$ such that
$\phi(r)=r^{\prime},\quad\tau^{\prime}\big{(}{\phi(a)}\big{)}=\phi\big{(}{\tau(a)}\big{)},\quad\delta^{\prime}(\phi(a))=\phi(\delta(a))+\phi(\tau(a))d-d\phi(a),$
@@
Proof. Let $\phi(y)=y^{\prime}+d^{\prime}$. Then $\phi$ extends from a map $\phi:A\rightarrow A^{\prime}$ to a homomorphism $\phi:A(\chi,r,\delta)\rightarrow A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$. We have
@@
$=\phi\big{(}\tau(a)\big{)}y^{\prime}+\phi\big{(}\tau(a)\big{)}d^{\prime}+\phi\big{(}\delta(a)\big{)}$
for any $a\in A$. Thus, we get
@@
Regarding the counit $\varepsilon$, we obtain
$d^{\prime}=(\varepsilon\otimes\iota)\big{(}\Delta(d^{\prime})\big{)}=\varepsilon(d^{\prime})+\varepsilon(r^{\prime})d^{\prime}=\varepsilon(d^{\prime})+d^{\prime},$
this indicates $\varepsilon(d^{\prime})=0$. By Proposition 3.3, we also know $\varepsilon(y)=0$ and $\varepsilon(y^{\prime})=0$. Hence,
@@
which shows that $\varepsilon\big{(}\phi(y)\big{)}=\varepsilon(y)$.
Now consider the antipode $S$, we find
$0=\varepsilon(d^{\prime})=m(S\otimes\iota)\Delta(d^{\prime})=S(d^{\prime})+S(r^{\prime})d^{\prime}=S(d^{\prime})+r{{}^{\prime-1}}d^{\prime},$
and therefore $S(d^{\prime})=-r{{}^{\prime-1}}d^{\prime}$. Consequently,
[…]
@@
Combining Theorem 3.9 and Theorem 4.2, we obtain the following corollary.
Corollary 4.3. Let $A$ and $A^{\prime}$ be multiplier Hopf *-coquasigroups, with their *-structures denoted by *, the corresponding Hopf-Ore extensions for $A$ and $A^{\prime}$ are $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$. If there exists a multiplier Hopf *-algebra isomorphism $\phi:A\rightarrow A^{\prime}$ such that the hypotheses of Theorem 4.2 are satisfied (where $(d^{\prime})^{*}=d^{\prime}$), and the following conditions hold:
(1) $\tau$ is an involution;
(2) $\delta(\tau+\iota)=0$ and $\delta*=*\delta$;
(3) $r^{*}=r$.
Then $A(\chi,r,\delta)$ and $A^{\prime}(\chi^{\prime},r^{\prime},\delta^{\prime})$ are isomorphic as Hopf-Ore extensions, and they are also isomorphic as multiplier Hopf *-coquasigroups.
Proof. By Theorem 3.9 and Theorem 4.2, it suffices to show that $\phi$ preserves the *-operation, that is, for any $x\in A(\chi,r,\delta)$, we have $\phi(x^{*})=\phi(x)^{*}$.
By the assumptions of this corollary, $\phi$ is a *-homomorphism on $A$, so for all $a\in A$, $\phi(a^{*})=\phi(a)^{*}$.
Moreover, by Theorem 4.2, we have $\phi(y)=y^{\prime}+d^{\prime}$. It follows that
$\phi(y^{*})=\phi(y)=y^{\prime}+d^{\prime},$
$\phi(y)^{*}=(y^{\prime}+d^{\prime})^{*}=y^{\prime}+(d^{\prime})^{*}=y^{\prime}+d,$
[…]